# Safe In-Context Reinforcement Learning

### v1

@@ Appendix E Ablation Studies
In this section, we present our ablation studies on reinforcement pretraining and supervised pretraining, examining shared factors such as context length and model size, as well as specific factors like dataset size for supervised pretraining. Our findings reveal distinct sensitivities to these factors. Reinforcement pretraining remains largely unaffected by model size, whereas supervised pretraining performance is significantly influenced by model size (See Figures 5.(b) and 6.(b)). Regarding context length, reinforcement pretraining benefits from longer sequences, while supervised pretraining exhibits degraded performance with longer contexts. Conversely, supervised pretraining performs better with shorter context lengths, where reinforcement pretraining struggles (See Figures 5.(a) and 6.(a)). These results indicate that learning long-term credit assignment is more challenging in offline reinforcement learning due to dataset constraints, whereas online learning, with access to environment interactions, manages long-term credit assignment more effectively. We also confirm the well-established finding that supervised pretraining is highly sensitive to dataset size, with models failing to learn and exhibiting random behavior on out-of-distribution data when the dataset is small (See Figure 6.(c)).
@@
Figure 5: Ablations of Safe Reinforcement Pretraining on SafeDarkRoom. (a), (b) The evaluation is set up similarly to Question (i). For context length, we compare 150 and 3000 against the base value of 1500. For model size, we compare embedding dimensions 32 and 128 against the base of 64. (c) At each training step, the average total return of 50 episodes across 10 random test environments, and the average Q Value across 50 episodes of 250 random train enironments are plotted. EPPO is easier to tune and more stable to train than the naive primal-optimal method.
@@
Figure 6: Ablations of Safe Supervised Pretraining on SafeDarkRoom. The evaluation is set up similarly to Question (i). For context length, we use a smaller base model and test three sequence lengths: 100, 500, and 1000. For dataset size, large refers to the full dataset, medium uses 50% of the dataset, and small uses 5% of the original dataset. Model size increases with the number of hidden layers: 2, 4, and 8, keeping other factors constant.

### v2

@@ Appendix E Ablation Studies
In this section, we present our ablation studies on SCARED and safe AD, examining shared factors such as context length and model size, as well as specific factors like dataset size for safe AD. Our findings reveal distinct sensitivities to these factors.
SCARED remains largely unaffected by model size, whereas safe AD performance is significantly influenced by model size (See Figures 5.(b) and 6.(b)). Regarding context length, SCARED benefits from longer sequences, while safe AD shows degraded performance with longer contexts. Conversely, safe AD performs better with shorter context lengths, where SCARED struggles (See Figures 5.(a) and 6.(a)). These results indicate that learning long-term credit assignment is more challenging in offline reinforcement learning due to dataset constraints, whereas online learning, with access to environment interactions, manages long-term credit assignment more effectively. We also confirm the well-established finding that safe AD is highly sensitive to dataset size, with models failing to learn and exhibiting random behavior on out-of-distribution data when the dataset is small (See Figure 6.(c)).
@@
Figure 5: Ablations of SCARED on SafeDarkRoom. (a), (b) The evaluation is set up similarly to Question (i). For context length, we compare 150 and 3000 against the base value of 1500. For model size, we compare embedding dimensions 32 and 128 against the base of 64. (c) At each training step, the average total return of 50 episodes across 10 random test environments, and the average Q Value across 50 episodes of 250 random train enironments are plotted. Our SCARED optimization is easier to tune and more stable to train than the naive primal-optimal method.
@@
Figure 6: Ablations of Safe Algorithm Distillation on SafeDarkRoom. The evaluation is set up similarly to Question (i). For context length, we use a smaller base model and test three sequence lengths: 100, 500, and 1000. For dataset size, large refers to the full dataset, medium uses 50% of the dataset, and small uses 5% of the original dataset. Model size increases with the number of hidden layers: 2, 4, and 8, keeping other factors constant.